Quantum Logistics, Inc., a wholesale distributor, is considering the construction of a new warehouse to serve the southeastern geographic region near the Alabama–Georgia border. There are three cities being considered. After site visits and a budget analysis, the expected income and costs associated with locating in each of the cities have been determined. The life of the warehouse is expected to be 12 years and MARR is 15%/year.

City

Initial Cost

Net Annual Income

Lagrange $1,260,000 $480,000
Auburn $1,000,000 $410,000
Anniston $1,620,000 $520,000

What is the future worth of each site?
Answer #1

t = 12 ys

MARR = 15%

FW of lagrange option = -1260000*(F/P,15%,12) + 480000*(F/A,15%,12)

= -1260000*5.350250 + 480000*29.001667

= 7179485.16

FW of auburn option = -1000000*(F/P,15%,12) + 410000*(F/A,15%,12)

= -1000000*5.350250 + 410000*29.001667

= 6540433.47

FW of aniston option = -1620000*(F/P,15%,12) + 520000*(F/A,15%,12)

= -1620000*5.350250 + 520000*29.001667

= 6413461.84